OpenStudy (anonymous):

find the vertex of the quadratic funciton f(x)=-2x^2+2x+3

OpenStudy (anonymous):

find the vertex, line of symmetry, the maximum or minumum value of the quadratic funciton f(x)=-2x^2+2x+3

OpenStudy (anonymous):

(1/2,9/2)

OpenStudy (anonymous):

the maximum is the vertex itself ,i.e.(1/2,9/2)

OpenStudy (anonymous):

line of symmetry is x=1/2

OpenStudy (anonymous):

by completing the square of the main formula (-1/2)(y-(11/6))=(x-(11/6))^2